# MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models

### v3

@@ Abstract
The tendency of Large Language Models (LLMs) to generate hallucinations raises concerns regarding their reliability. Therefore, confidence estimations indicating the extent of trustworthiness of the generations become essential. However, current LLM confidence estimations in languages other than English remain underexplored. This paper addresses this gap by introducing a comprehensive investigation of Multilingual Confidence estimation (MlingConf) on LLMs, focusing on both language-agnostic (LA) and language-specific (LS) tasks to explore the performance and language dominance effects of multilingual confidence estimations on different tasks. The benchmark comprises four meticulously checked and human-evaluate high-quality multilingual datasets for LA tasks and one for the LS task tailored to specific social, cultural, and geographical contexts of a language. Our experiments reveal that on LA tasks English exhibits notable linguistic dominance in confidence estimations than other languages, while on LS tasks, using question-related language to prompt LLMs demonstrates better linguistic dominance in multilingual confidence estimations. The phenomena inspire a simple yet effective native-tone prompting strategy by employing language-specific prompts for LS tasks, effectively improving LLMs’ reliability and accuracy on LS tasks.
@@
Boyang Xuethanks:   Equal Contribution. , Hongru Wang♠∗, Rui Wang, Sheng Wang Zezhong Wang, Yiming Du, Bin Liang, Kam-Fai Wong The Chinese University of Hong Kong The University of Hong Kong {byxue, hrwang, kfwong}@se.cuhk.edu.hk
@@ 1 Introduction
LLMs’ susceptibility to generating hallucinated contents in their generations incurs concerns about unreliability in real-world applications Ji et al. (2023); Rawte et al. (2023). Therefore, it becomes increasingly crucial for users to directly ascertain how much they can trust a model’s response. Assessing the confidence or uncertainty of a model’s output can immediately indicate the level of reliability to users, thereby playing a key role in developing trustworthy AI systems Geng et al. (2023); Kadavath et al. (2022).
@@
However, existing research on confidence or uncertainty estimations of LLMs has been confined exclusively in English Kadavath et al. (2022); Lin et al. (2022); Geng et al. (2023); Tian et al. (2023b). Considering the LLMs’ global deployment, the dearth of confidence estimations in languages other than English hinders users from ascertaining the reliability of generations in non-English scenarios. Owing to the disparities in the quantity and domain coverage of pre-training corpora across different languages, the confidence estimation ability may also presumably vary. The performance of confidence estimation methods mainly employed in English remains a subject worthy of exploration when generalized to other languages.
@@
Additionally, linguistic dominance referring to that one language holds a superior status over others within a specific social or cultural context (Blommaert, 2010; Treffers-Daller, 2019; Heller, 2007), also exist in confidence estimation ability on different languages. To achieve more reliable confidence estimations in multilingual contexts over various tasks, we study and divide the tasks into language-agnostic (LA) (Zhao et al., 2020) and language-specific (LS) scenarios as in Figure 1. In this study, the LS refers to the tasks that hold linguistic dominance caused by the knowledge domain coverage varying in different language training corpora, such as questions pertaining to social, cultural, or geographical contexts for a specific language, while the LA task involves linguistic dominance mainly caused by the quantities of training corpora, such as general knowledge, common sense, and reasoning (Basaj et al., 2018; Sánchez et al., 2024).
@@
To this end, we propose a benchmark called MlingConf (Multilingual Confidence) to investigate the performance of several LLM confidence estimation methods on five languages including English, Japanese, Chinese, French, and Thai on both LA and LS tasks respectively. First, we meticulously construct the MlingConf dataset for this study including five high-quality multilingual datasets in LA and LS scenarios. The LA involves four tasks including closed-book factual knowledge question-answering (QA) derived from TriviaQA Joshi et al. (2017), math reasoning from GSM8K (Cobbe et al., 2021), common sense reasoning from CommonsenseQA (Talmor et al., 2019), and professional scientific knowledge QAs from (Johannes Welbl, 2017). These four datasets widely used in confidence estimation in English (Kuhn et al., 2023; Xiong et al., 2024) are translated into other four languages. We also create a language-specific QA (LSQA) for the LS scenario, which is comprised of specific social, cultural, and geographical QA pairs for the investigated five languages respectively. To ensure the data quality, we conduct rigorous translation consistency checks to filter the failed samples and finally employ linguistic experts for human evaluation.
@@
Experiments are conducted on three major LLM confidence estimation methods including probability-based (Vazhentsev et al., 2023; Varshney et al., 2023), $p(\mathrm{True})$-based (Kadavath et al., 2022), and self-verbalized (Lin et al., 2022; Xiong et al., 2024) confidence estimations using the curated five multilingual datasets on several LLMs. We evaluate the confidence estimation ability and calibration using AUROC and ECE. Results on LA tasks suggest that all the confidence estimation methods demonstrate comparable performance of reliability across different languages, while the calibration performance in English significantly outperforms others. Results on the LS task reveal a pronounced phenomenon of language dominance, indicating that, for questions related to specific linguistic contexts, utilizing the respective languages yields the highest accuracy and confidence estimation performance. This observation inspires a native-tone prompting strategy, whereby, in the LS task, the relevant linguistic background of the question is first assessed, and then the corresponding language is employed for the response. Compared to the use of any single language, this approach leads to significant improvements in both accuracy and confidence estimation.
@@
$\bullet$ To the best of our knowledge, the MlingConf first propose to investigate multilingual confidence estimations with intricately constructed and expert-checked MlingConf datasets for both LA and LS scenarios, serving as a valuable benchmark for future works  [^1] The codes and the MlingConf datasets have been released on https://github.com/AmourWaltz/MlingConf.
$\bullet$ We conduct experiments on MlingConf datasets and the results reveal that language dominance affects confidence estimation performance in English on LA tasks and query-related languages on LS tasks respectively.
$\bullet$ We propose a native-tone prompting strategy, which significantly enhances the reliability and accuracy compared to the use of any single language prompts for LS tasks.
@@ 2.1 Data Source
Owing to the lack of multilingual resources to comprehensively exhibit confidence estimation across diverse languages, our study leverages several datasets in English (en) and translates into four languages, encompassing Japanese (ja), Chinese (zh), French (fr), and Thai (th). We evaluate the quality of confidence estimates on our proposed MlingConf dataset including four datasets for the language-agnostic (LA) scenario and one dataset for the language-specific (LS) scenario.
@@
For LA tasks, we employ the following four datasets that are widely used for confidence estimation research in English in Kuhn et al. (2023); Xiong et al. (2024). We translate the datasets into other four languages. The datasets include 1) TriviaQA (TVQA) (Joshi et al., 2017) of closed-book trivia question-answering pairs to gauge models’ factual knowledge; 2) GSM8K (Cobbe et al., 2021) for arithmetic reasoning task of math problems; 3) CommonsenseQA (CSQA) (Talmor et al., 2019) of multiple-choice QA pairs requiring different types of commonsense knowledge; 4) SciQ (Johannes Welbl, 2017) requiring scientific professional knowledge.
@@
We create Language-Specific QA (LSQA) dataset pertaining to language-dominant knowledge covering specific social, geographical, and cultural language contexts for the UK & US, France, China, Japan, and Thailand respectively. We prompt GPT-4 OpenAI (2023)  [^2] https://platform.openai.com/docs/api-reference to generate 200 questions pertaining to only one specific language contexts as a language-specific subset. As demonstrated in Figure 12, for example, all questions in Japanese subset pertain to Japanese social culture, history, or geography. All the subsets are also translated into other four languages.
The language selection, dataset details, prompts, and samples extracted from the LSQA are demonstrated in Appendix A, B, and C respectively.
Figure 2: Three stages of MlingConf dataset construction.
@@ 2.2 Dataset Construction
The construction of the MlingConf dataset in this study follows a three-stage procedure as delineated in Figure 2.
@@
: The QA samples derived from the above datasets are translated into four languages (ja, th, zh, fr) through GPT-4 OpenAI (2023);
@@
: We check the consistency of five translated results by comparing the semantic equivalence in pairs in $C_{5}^{2}=10$ times for each sample. The samples with more than 2 times semantic in-equivalence are treated as noisy data and then filtered. The changes of samples before and after consistency check and filter are shown in Table 1 and more clean multilingual datasets are obtained. Moreover, we present the number of samples for each language-specific subset in the LSQA dataset as in Table 2.
@@
Stage 3
: Finally, we employ several experts majoring in linguistics to examine the translation performance across 50 randomly selected samples as shown in Table 3. Human evaluation results suggest the obtained multilingual datasets are high-quality for further experiments. Temperature $T$ is set to 0 for all the generations. The translation and semantic equivalence comparison prompts are presented in Appendix C.
Lang. TVQA GSM8K CSQA SciQ LSQA
zh 96% 100% 100% 98% 100%
ja 98% 100% 98% 96% 100%
fr 100% 100% 100% 98% 100%
th 96% 100% 98% 94% 100%
@@ 3.1 Confidence Estimation Methods
The $p(\mathrm{True})$ confidence score is implemented by simply asking the model itself if its first proposed answer $\boldsymbol{y}$ to the question $\boldsymbol{x}$ is true Kadavath et al. (2022), and then obtaining the probability $p(\mathrm{True})$, which can implicitly reflect self-reflected certainty.
@@ 3.2 Evaluation Metrics
AUROC assesses the effectiveness of confidence estimation Filos et al. (2019) by quantifying how likely a randomly chosen correct answer possesses a higher confidence score than an incorrect one, yielding a score within the range of [0, 1], implemented by sklearn toolkit  [^3] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py. A higher AUROC score implying higher reliability is preferred.
@@
3.3 Implementation Details
The experiments are conducted on three LLMs: GPT-3.5-Turbo-0125 (GPT-3.5), Llama-3.1-8B-Instruct (Llama-3.1)  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024), and Llama-2-13B-Chat (Llama-2)  [^5] https://huggingface.co/meta-llama/Llama-2-13b-chat (Touvron et al., 2023), and Vicuna-7B-v1.5 (Vicuna-v1.5)  [^6] https://huggingface.co/lmsys/vicuna-7b-v1.5 (Zheng et al., 2023). We only present the results of the current most commonly used commercial LLM GPT-3.5 and open-source LLM Llama-3.1 in the main part and leave others in Appendix E. Few-shot prompts containing $N_{f}$ examples are utilized for answer generation with the greedy decoding where $N_{f}$ is set to 8 for GSM8K and 4 for others.
4 Experimental Results and Analysis
Dataset Conf. en zh ja fr th Avg.
@@ 4 Experimental Results and Analysis
CSQA Prob. 59.06 24.45 55.92 38.30 48.01 50.60 55.33 31.12 48.21 41.71 53.31 48.18
@@
SciQ Prob. 69.50 32.23 71.29 35.63 78.28 47.06 72.66 34.85 75.13 56.17 73.37 41.19
@@
Avg. Prob. 64.96 26.88 66.02 33.51 65.07 38.71 64.37 32.66 64.88 44.95 65.06 35.34
@@
Llama-3.1
TVQA Prob. 80.74 10.72 80.41 40.39 88.75 26.24 79.05 20.38 89.59 36.77 83.71 26.90
GSM8K Prob. 65.69 22.33 66.37 21.92 69.73 35.69 61.07 29.56 63.22 28.51 65.22 27.60
CSQA Prob. 78.06 13.64 64.91 36.33 75.65 30.11 66.12 19.72 77.65 42.18 72.47 28.40
SciQ Prob. 74.14 13.40 72.09 32.26 74.48 34.21 77.45 22.76 77.61 36.10 75.15 27.75
Avg. Prob. 74.88 15.02 70.94 32.72 77.15 31.56 70.92 23.10 77.01 35.89 74.14 27.66
@@
In the experiment part, we aim to address the following three research questions:
a. How do performance (ARC. and ECE) vary in three confidence estimation methods on LA tasks?
b. How do performance and accuracy vary in different languages on LA tasks?
c. How do performance and accuracy vary in language usage on different language-specific LSQA subsets on LS tasks?
4.1 Results and Analysis on LA Tasks
We present the results of the confidence estimation performance of four LA datasets using GPT-3.5 and Llama-3.1 as presented in Table 4. The key findings and analysis for research questions a and b are listed below.
@@ 4.1 Results and Analysis on LA Tasks
Recalling research question a, we highlight the supreme performance using   among the three methods for each column of the dataset in Table 4. On GPT-3.5, both p(True) and Verb. yield the superior performance than Prob. across all languages averaged on four datasets (ARC.: 70.65, 70.45 vs. 66.02; ECE: 29.66, 28.64 vs. 35.34). p(True) and Verb. have comparable performance in ARC. scores, while Verb. is better calibrated. In contrast, Prob. show the superior performance than p(True) and Verb. and performs more stable on Llama-3.1 (ARC.: 74.14 vs. 61.51 and 62.92; ECE: 27.66 vs. 26.54 and 35.93). p(True) demonstrates better calibration results on languages other than English.
@@
The performance differences between GPT-3.5 and Llama-3.1 can be attributed to that GPT-3.5’s strong instruction-following abilities benefits the prompt-based confidence estimation methods Verb. and p(True)., while Llama-3.1’s likelihood probabilities are well-calibrated during training.
Recalling research question b, previous studies only validate the efficacy of prompt-based confidence assessments exclusively within English. Our findings indicate that the approaches are also preferable in other languages and the performance fluctuates in different languages. Specifically as demonstrated in Table 4, ARC. scores are less fluctuating across different languages while ECE in English (23.04 and 22.77) performs better than in other languages on both GPT-3.5 and Llama-3.1.

### v4

@@ Abstract
The tendency of Large Language Models (LLMs) to generate hallucinations raises concerns regarding their reliability. Therefore, confidence estimations indicating the extent of trustworthiness of the generations become essential. However, current LLM confidence estimations in languages other than English remain underexplored. This paper addresses this gap by introducing a comprehensive investigation of Multilingual Confidence estimation (MlingConf) on LLMs, focusing on both language-agnostic (LA) and language-specific (LS) tasks to explore the performance and language dominance effects of multilingual confidence estimations on different tasks. The benchmark comprises four meticulously checked and human-evaluated high-quality multilingual datasets for LA tasks and one for the LS task tailored to specific social, cultural, and geographical contexts of a language. Our experiments reveal that on LA tasks English exhibits notable linguistic dominance in confidence estimations than other languages, while on LS tasks, using question-related language to prompt LLMs demonstrates better linguistic dominance in multilingual confidence estimations. The phenomena inspire a simple yet effective native-tone prompting strategy by employing language-specific prompts for LS tasks, effectively improving LLMs’ reliability and accuracy in LS scenarios.
@@
[…]
@@ 1 Introduction
Large Language Models’ (LLMs) susceptibility to generating hallucinated contents incurs concerns about unreliability in real-world applications Ji et al. (2023); Rawte et al. (2023). Therefore, it becomes increasingly crucial for users to directly ascertain how much they can trust a model’s response. Assessing the confidence or uncertainty of a model’s output can immediately indicate the level of reliability to users, thereby playing a key role in developing trustworthy AI systems Geng et al. (2023); Kadavath et al. (2022).
@@
However, existing research on confidence or uncertainty estimations for LLMs has been predominantly limited to English Kadavath et al. (2022); Lin et al. (2022); Geng et al. (2023); Tian et al. (2023b). The dearth of confidence estimations in languages other than English hinders users from assessing the reliability of LLMs in non-English scenarios, restricting the LLMs’ global deployment. Due to the variations in the quantity and domain coverage of pre-training corpora across different languages, the confidence estimation ability may also presumably vary. Therefore, the performance of confidence estimation methods primarily developed for English remains a crucial subject for explorations when applied to other languages.
@@
Additionally, to conduct a fine-grained investigation of multilingual confidence estimations across various tasks, we divide the tasks into language-agnostic (LA) (Zhao et al., 2020) and language-specific (LS) scenarios as in Figure 1 considering the effects of linguistic dominance. Linguistic dominance refers to that one language holds a superior status over others within a specific social or cultural context (Blommaert, 2010; Treffers-Daller, 2019; Heller, 2007), can also exist in confidence estimation ability on different languages. In this study, the LS refers to the tasks that hold linguistic dominance caused by the knowledge domain coverage varying in different language training corpora, such as questions pertaining to social, cultural, or geographical contexts for a specific language, while the LA involves linguistic dominance mainly caused by the quantities of training corpora, such as general knowledge, common sense, and reasoning (Basaj et al., 2018; Sánchez et al., 2024).
@@
To this end, we propose a benchmark called MlingConf (Multilingual Confidence) to investigate the performance of several LLM confidence estimation methods on five languages including English, Japanese, Chinese, French, and Thai. First, we meticulously constructed a high-quality multilingual dataset called the MlingConf dataset for the benchmark including five datasets of different tasks in LA and LS scenarios respectively. The LA involves four different tasks that are widely used in confidence estimation in English (Kuhn et al., 2023; Xiong et al., 2024) are translated into other four languages. We also create a language-specific QA (LSQA) dataset for the LS scenario, including five subsets for the investigated five languages respectively. Each subset comprises QA pairs about social culture, history, geography, and celebrity pertaining to the specific language. To ensure the data quality, we conduct rigorous translation consistency checks to filter the failed samples and finally employ linguistic experts for human evaluations.
@@
Experiments are conducted on three major LLM confidence estimation methods including probability-based (Vazhentsev et al., 2023; Varshney et al., 2023) and prompt-based confidence estimations ($p(\mathrm{True})$ (Kadavath et al., 2022) and self-verbalize (Lin et al., 2022; Xiong et al., 2024)) using the curated five multilingual datasets on several LLMs. We evaluate the confidence estimation ability and calibration using AUROC and ECE. Results on LA tasks suggest that prompt-based confidence estimations are preferable on LLMs with stronger instruction-following abilities, and English exhibits linguistic dominance. Results on the LS task reveal a pronounced phenomenon of language dominance, indicating that, for questions related to specific linguistic contexts, utilizing the respective languages yields the highest accuracy and confidence estimation performance. This observation inspires a native-tone prompting strategy: whereby, in the LS task, the relevant linguistic background of the question is first assessed, and then the corresponding language is employed to generate the response. Compared to the use of any single language, this approach leads to significant improvements in both accuracy and confidence estimations. Furthermore, we employ and generalize on extended confidence estimation methods and languages for both LS and LA tasks. The results further complete and enhance the above findings and analysis to the benchmark.
@@
$\bullet$ To the best of our knowledge, the MlingConf first proposes to investigate multilingual confidence estimations with intricately constructed and expert-checked MlingConf datasets for both LA and LS scenarios, serving as a valuable benchmark for future works of reliable multilingual LLMs  [^1] The codes and the MlingConf datasets have been released on https://github.com/AmourWaltz/MlingConf..
$\bullet$ Experiments conducted on MlingConf datasets present valuable findings about confidence estimation uses in multilingual scenarios, language dominance effects of English on LA tasks, and query-related languages on LS tasks respectively.
$\bullet$ Based on the observed linguistic dominance on LS tasks, we propose a native-tone prompting strategy, which significantly enhances the reliability and accuracy compared to the use of any single language prompts for LS tasks.
@@ 2 MlingConf Dataset
Owing to the lack of multilingual resources to comprehensively exhibit confidence estimation across diverse languages, we construct a high-quality multilingual dataset called MlingConf dataset encompassing five languages: English (en), Japanese (ja), Chinese (zh), French (fr), and Thai (th). Specifications of the language selection in consideration of language family and resource level are demonstrated in Appendix A. The MlingConf dataset includes four tasks for the language-agnostic (LA) scenario and one task for the language-specific (LS) scenario. We specify the data source and construction process of the MlingConf dataset in Sec. 2.1 and 2.2 respectively. Further dataset details and statistics can be referred to Appendix B.
Figure 2: Three stages of MlingConf dataset construction.
@@ 2.1 Data Source
For LA tasks, we employ the following four datasets that are widely used for confidence estimations in English (Kuhn et al., 2023; Xiong et al., 2024). The datasets include 1) TriviaQA (TVQA) (Joshi et al., 2017) of closed-book trivia question-answering pairs to gauge models’ factual knowledge; 2) GSM8K (Cobbe et al., 2021) for arithmetic reasoning task of math problems; 3) CommonsenseQA (CSQA) (Talmor et al., 2019) of multiple-choice QA pairs requiring different types of commonsense knowledge; 4) SciQ (Johannes Welbl, 2017) requiring scientific professional knowledge. All the datasets are pre-processed in standard QA format.
@@
We create Language-Specific QA (LSQA) dataset pertaining to language-dominant knowledge covering specific social, geographical, and cultural language contexts for the UK & US, France, China, Japan, and Thailand respectively. We prompt GPT-4 OpenAI (2023)  [^2] https://platform.openai.com/docs/api-reference to generate 200 questions pertaining to only one specific language contexts as a language-specific subset. As demonstrated in Figure 11, for example, all questions in Japanese subset pertain to Japanese social culture, history, geography, celebrities, etc.
@@ 2.2 Dataset Construction
The construction of the MlingConf dataset in this study follows an elaborate three-stage procedure as delineated in Figure 2.
@@
The QA samples derived from the above datasets are translated into four languages (ja, th, zh, and fr) through GPT-4 OpenAI (2023).
@@
We check the consistency of five translated results by comparing the semantic equivalence in pairs in $C_{5}^{2}=10$ times for each sample. The samples with more than 2 times semantic in-equivalence are treated as noisy data and then filtered. The changes of samples before and after consistency check and filter are in Table 2 and more clean multilingual datasets are obtained. Moreover, we present the number of samples for each language-specific LSQA subset as in Table 3.
Stage 3
Finally, we employ several experts majoring in linguistics to examine the translation performance across 50 randomly selected samples as shown in Table 1. Given the dataset obtained after Stage 2, we randomly select 50 samples from each test set and send them to language experts. Each language expert is required to evaluate the translation results of their respective specialized language in each sample, determining whether the translation is correct (returning 1 for correct and 0 otherwise). We calculate the translation accuracy for each language in each test set and present the results in Table 1. Human evaluation results suggest the obtained multilingual datasets are high-quality for further experiments.
For all generations of MlingConf dataset construction, the temperature $T$ is set to 0. The translation and semantic equivalence comparison prompts are presented in Appendix C.
Lang. TVQA GSM8K CSQA SciQ LSQA
zh 96% 100% 100% 98% 100%
ja 98% 100% 98% 96% 100%
fr 100% 100% 100% 98% 100%
th 96% 100% 98% 94% 100%
@@ 3.1 Confidence Estimation Methods
The $p(\mathrm{\mathrm{True}})$ confidence score is implemented by simply asking the model itself if its first proposed answer $\boldsymbol{y}$ to the question $\boldsymbol{x}$ is true Kadavath et al. (2022), and then obtaining the probability $p(\mathrm{True})$, which can implicitly reflect self-reflected certainty.
@@ 3.2 Evaluation Metrics
AUROC assesses the effectiveness of confidence estimation Filos et al. (2019) by quantifying how likely a randomly chosen correct answer possesses a higher confidence score than an incorrect one, yielding a score in range of [0, 1], implemented by sklearn toolkit  [^3] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py.
@@
Conf. en zh ja fr th Avg.
@@
GSM8K on Llama-3.1
Prob. 65.69 22.33 66.37 21.92 69.73 35.69 61.07 29.56 63.22 28.51 65.22 27.60
CSQA on GPT-3.5
Prob. 59.06 24.45 55.92 38.30 48.01 50.60 55.33 31.12 48.21 41.71 53.31 48.18
@@
CSQA on Llama-3.1
Prob. 78.06 13.64 64.91 36.33 75.65 30.11 66.12 19.72 77.65 42.18 72.47 28.40
SciQ on GPT-3.5
Prob. 69.50 32.23 71.29 35.63 78.28 47.06 72.66 34.85 75.13 56.17 73.37 41.19
@@
SciQ on Llama-3.1
Prob. 74.14 13.40 72.09 32.26 74.48 34.21 77.45 22.76 77.61 36.10 75.15 27.75
Avg. (TVQA, GSM8K, CSQA, SciQ) on GPT-3.5
Prob. 64.96 26.88 66.02 33.51 65.07 38.71 64.37 32.66 64.88 44.95 65.06 35.34
@@
Avg. (TVQA, GSM8K, CSQA, SciQ) on Llama-3.1
Prob. 74.88 15.02 70.94 32.72 77.15 31.56 70.92 23.10 77.01 35.89 74.14 27.66
@@
3.3 Implementation Details
Experiments are conducted on GPT-3.5-Turbo-0125 (GPT-3.5) and Llama-3.1-8B-Instruct (Llama-3.1)  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct (AI@Meta, 2024). We only present the results of the current most commonly used commercial GPT-3.5 and open-source Llama-3.1 in the main part and leave the results on some other LLMs in Appendix E. Few-shot prompts containing $N_{f}$ examples are utilized for answer generation with greedy decoding which outperforms temperature decoding on knowledge tasks (Song et al., 2024). $N_{f}$ is set to 8 for GSM8K and 4 for others.
4 Experiments on LA Tasks
To comprehensively investigate LLMs’ multilingual confidence estimations on LA tasks, as presented in Table 4 and Figure 3, experiments are conducted to observe performances varying in different confidence estimation methods and languages in Sec. 4.1 and 4.2 respectively.
@@ 4 Experiments on LA Tasks
4.1 Results regarding Different Confidence Estimations on LA Tasks
Findings:
@@ 4.1 Results regarding Different Confidence Estimations on LA Tasks
Applying prompt-based confidence estimations is preferable in multilingual tasks for LLM with stronger instruction-following ability. The probability method performs better confidence estimations on relatively weak LLM. The findings provide a direct takeaway about selecting optimal confidence estimations of LLMs in multilingual scenarios. In Table 4, we highlight the supreme performance in bold among the three methods for each column of each dataset. On GPT-3.5, both p(True) and Verb. yield the superior performances than Prob. across all languages averaged on four datasets (ARC.: 70.65, 70.45 vs. 66.02; ECE: 29.66, 28.64 vs. 35.34). p(True) and Verb. have comparable performance in ARC. scores, while Verb. is better calibrated. In contrast, Prob. shows superior performance than p(True) and Verb. and performs more stable on Llama-3.1 (ARC.: 74.14 vs. 61.51 and 62.92; ECE: 27.66 vs. 26.54 and 35.93). p(True) demonstrates better calibration results on languages other than English.
Analysis:
@@
In Table 4, the performance differences between two LLMs can be attributed to that GPT-3.5’s strong instruction-following abilities benefit the prompt-based multilingual confidence estimation methods Verb. and p(True), but leading to over-confidence in output token probabilities. In contrast, Llama-3.1 cannot stably generate verbalized confidence scores and perform relatively weak instruction-following abilities, but maintain well-calibrated likelihood probabilities during the pre-training stage for all languages.
4.2 Results regarding Different Languages on LA Tasks
Findings:
Linguistic dominance is manifested in English with superior confidence estimation performances on LA tasks for multilingual LLMs. Prior works only validate the efficacy of prompt-based confidence estimations in English. Our findings indicate that the methods are also preferable in other languages and performances fluctuate in different languages. In Table 4, ARC. scores are less fluctuating across different languages while ECE in English (23.04 and 22.77) performs better than in other languages on both GPT-3.5 and Llama-3.1. We also report the accuracy on LA datasets in Figure 3. English consistently performs better across all datasets exclusively GSM8K. Generally, prompting in English outperforms others, hence responding in English on LA tasks can be adequately credible and accurate where linguistic dominance is leading in English.
Analysis: